OpenStudy (briana.img):

The perimeter of a rectangular field is 200 feet. The length of the field is 7 feet more than two times the width of the field.

OpenStudy (briana.img):

Part A Write a system of two equations that could be used to find the length and width of the field. Let x represent the length of the field, and let y represent the width of the field. Part B Solve the system of equations for x and y to determine the dimensions of the field.

OpenStudy (anonymous):

the first one should be ok right \[2x+2y=200\] or maybe even \[x+y=100\]

OpenStudy (briana.img):

so just the substitution method?

OpenStudy (anonymous):

hold on you still need a second equation

OpenStudy (briana.img):

don't they both work as two?!

OpenStudy (anonymous):

The length of the field is 7 feet more than two times the width of the field

OpenStudy (briana.img):

oh nvm

OpenStudy (anonymous):

oh i have confused you the perimeter is \(2x+2y=200\)i just made it easier and wrote \(x+y=100\) it is the same equation

OpenStudy (briana.img):

oooh ok

OpenStudy (anonymous):

how about translating The length of the field is 7 feet more than two times the width of the field

OpenStudy (briana.img):

addition

OpenStudy (briana.img):

7ft + 2x = width of field ??

OpenStudy (anonymous):

yeah i would say \[x=2y+7\] i think

OpenStudy (mathmale):

If i were you, I'd take the information contained in the problem statement, "The perimeter of a rectangular field is 200 feet. The length of the field is 7 feet more than two times the width of the field," choose variables to represent the length and width, and then write two equations that accurately represent the verbal statements. Only then should you consider how to solve this system of linear equations.

OpenStudy (anonymous):

it really doesn't matter but they told you \(x\) is the length and \(y\) is the width they are being annoying and telling you how to do it as if there is one method

OpenStudy (briana.img):

@satellite73 tru tru

OpenStudy (anonymous):

now that you have your two equations \[x+y=100\\ x=2y+7\] you can solve

OpenStudy (briana.img):

2x(7y+7)+2y=200 ?

OpenStudy (briana.img):

accidentally added the x

OpenStudy (anonymous):

yeah try again i think you will get it

OpenStudy (briana.img):

2(7y+7)+2y=200 14y+14+2y=200

OpenStudy (briana.img):

14+16y=200

OpenStudy (anonymous):

i think one of those sevens should be a two

OpenStudy (anonymous):

not \[2(7y+7)+2y=200 \] but rather \[2(2y+7)+2y=200 \]

OpenStudy (briana.img):

4y+14+2y=200 14+6y=200 6y=200-14=186 ?

OpenStudy (briana.img):

6y=186

OpenStudy (anonymous):

ok

OpenStudy (briana.img):

y=31

OpenStudy (anonymous):

looks good to me

OpenStudy (briana.img):

x+31=186 x=155

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\[2x+2y=200\\ 2x+2\times 31=200\]

OpenStudy (anonymous):

it isn't \(x+y=200\) it is either \[2x+2y=200\] or \[x+y=100\]

OpenStudy (briana.img):

awh the answers i came up with are wrong

OpenStudy (anonymous):

really?

OpenStudy (briana.img):

yeah 155+31=186 2(155)+2(31)=372

OpenStudy (anonymous):

i get \[x=69,y=31\]

OpenStudy (anonymous):

wait i see the problem you had \(x+31=200\) but is should have been \(x+31=100\)

OpenStudy (briana.img):

ooh

OpenStudy (briana.img):

where exactly was where i messed?

OpenStudy (briana.img):

i dont see it

OpenStudy (anonymous):

you wrote \[x+31=186\] i have no idea where that came from

OpenStudy (anonymous):

it could have been either \[x+31=100\] or \[2x+2\times 31=200\] or easier still \[x=2\times 31+7\]

OpenStudy (briana.img):

OOOH okay i see it now

OpenStudy (briana.img):

thank you so much for being patient with me!!

OpenStudy (anonymous):

yw good luck!